Match the following aqueous solutions with the appropriate letter from the column on the right.

1. 0.16 m NiBr2 A. Lowest freezing point
2. 0.15 m Mg(NO3)2 B. Second lowest freezing point
3. 0.20 m MnSO4 C. Third lowest freezing point
4. 0.42 m Sucrose(nonelectrolyte) D. Highest freezing point

2.The boiling point of diethyl ether, CH3CH2OCH2CH3, is 34.500 °C at 1 atmosphere. Kb(diethyl ether) = 2.02 °C/m

In a laboratory experiment, students synthesized a new compound and found that when 10.97 grams of the compound were dissolved in 291.1 grams of diethyl ether, the solution began to boil at 34.779 °C. The compound was also found to be nonvolatile and a non-electrolyte.


What is the molecular weight they determined for this compound ?

g/mol
Answer #1

1) The influence of the molal concentration in the ionic compounds, grows proportionally to the amount of ions in which it dissociates, the higher the molality, the lower the freezing point, then the match is made:

1 - A

2 - B

3 - D

4 - C

2) The molality of the solution is calculated:

m = ΔT / Kb = (34.779 - 34.5) / 2.02 = 0.138 m

The moles of the solute are calculated:

n = m * Kg solvent = 0.138 m * 0.2911 Kg = 0.040 mol

The molar weight is calculated:

MM = m / n = 10.97 g / 0.04 mol = 274.25 g / mol

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.